OpenStudy (goformit100):

Today is Varun's birthday. One year, from today he will be twice as old as he was 12 years ago. How old is Varun today ? A. 20 years B. 22 years C. 25 years D. 27 years

OpenStudy (goformit100):

@Eyad

OpenStudy (anonymous):

hmmm , Let Varan's age today = x years. Then, Varan's age after 1 year = (x + 1) years. Therefore x + 1 = 2 (x - 12) x + 1 = 2x - 24 x = 25 ====== Easy right ?

OpenStudy (goformit100):

Yes Baby. :) ooops sorry Mr. @Eyad